# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

### v2

@@ 5.1 Baselines for comparison
  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method that relies on using a heuristic-based strategy for retaining layers at fixed intervals. We replicate the algorithm in (Liu et al., 2024) and compare performance both with and without LoRA fine-tuning across various skipping percentages.
@@ 5.2 Experimental Results on Different Tasks
WMT: Our experiments demonstrate that FiRST performs consistently well in different machine translation tasks (EN-to-DE, EN-to-ZH) across different models. For LLaMA-3.8B, for $15\%$ skipping, FiRST achieves a latency improvement of 10-12% on TPOT (see Tables 1 and 5) while being competitive to the base model + LoRA (gold reference) in quality. In most cases, it significantly outperforms other layer skipping strategies (Random, Skip Decode, Unified Skipping) and in other cases, it is comparable in quality. In comparison to the gold output (Base + LoRA), FiRST is $\geq 80\%$ in COMET scores (semantic metric), $\geq 85\%$ in BLEU-1 scores, and $\geq 75\%$ in BLEU-2 scores (syntax metrics) for both EN-to-DE and EN-to-ZH translations. For $25\%$ skipping, FiRST achieves significant improvement in quality over other strategies, in almost all metrics, while achieving $\sim 18\%$ reduction in TPOT.
For LLaMA-3.2-3B, the latency improvement is $\sim 10\%$, with quality scores being significantly higher than other layer skipping strategies (Table 6). It is within $65-85\%$ of BLEU-1 and COMET scores (Table 2) for EN-DE and EN-ZH of the gold (LoRA fine-tuned base model).
@@
CNN/DailyMail Dataset: For LLaMA-3-8B, at roughly 15% skipping level, our method outperforms the base model + LoRA (Table: 4) while obtaining a 12% improvement in TPOT (Table 5). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.88x
R + LoRA 25 0.82x 0.83x
R + LoRA 35 0.69x 0.68x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 20 0.81x
R + LoRA 27 0.76x
@@
Table 5: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.91x
R + LoRA 25 0.78x 0.75x
R + LoRA 35 0.69x 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 24 0.79x
R + LoRA 28 0.77x
Table 6: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.83x
R + LoRA 28 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.84x
R + LoRA 28 0.72x
Table 7: TPOT variation of LLaMA-3-8B (left) and LLaMA-3.2-3B (right) on SQuAD dataset for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
@@
SQuAD Dataset: For the LLaMA-3-8B model, FiRST is $>95\%$ in output quality of gold (base + LoRA) (Table 3), with overall latency gains of 6-16% (Table 7). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 3) with gains in latency of 6-16% overall (Table 7) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics. Detailed results for an additional skipping percentage are provided in Appendix A.8.
@@
Layer-wise Skipping Patterns: Layer-wise skipping varies significantly across tasks, reflecting the task-specific importance of each layer. For LLaMA-3-8B at a 15% skipping rate, layers 7–9 and 21 are fully skipped in English-to-German, with partial skipping in layer 18. In English-to-Chinese, layers 7–9, 16, and 21 are fully skipped, while layer 20 is partially skipped. For Summarization, layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19, 21, and 26. Some layers are skipped less than 10%, indicating their necessity for specific sequences. This task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped, and skipping patterns depend on the input. Detailed statistics are in Appendix A.7.
Computational overhead of Routers: We also examine the computational overhead introduced by training the routers. This overhead remains minimal, as the router parameters are significantly smaller than the total number of trainable parameters in each model (0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B). The proportion of the total computational time spent on the router operations compared to the entire forward pass of the model is close to 0.3% for both models suggesting their overhead does not significantly affect the efficiency of the model during inference or training.
@@ 6 Conclusion
We propose a new framework FiRST for layer selection corresponding to input sequence and task towards reducing latency in a quality aware manner. This is sequence dependent and operates in a KV cache compatible manner. For an optimal skipping rate of around 15%, FiRST achieves 10-20% reduction in latency while being quality neutral (approximately 80% or more in quality metrics compared to base model) on multiple tasks/model architectures such as Machine Translation, Summarization and Question Answering, on well known open source datasets, and some cases, even improving upon base model, while significantly outperfoming other layer selection strategies on most quality metrics.
@@ A.1 Related Work
Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022) where adaptive compute is used for different parts of the token sequence. While these approaches have been popular for encoder-only models which processes the entire sequence as a whole, they have faced challenges in generation tasks. The main limitation of these set of techniques are their inability to handle KV caching appropriately which is crucial for multi-fold speed-ups in current LLM architectures. We emphasize that in our work, we assign varying compute to sequences in different batches but within the same sequence, we assign the same compute to every token.
@@ A.2 Details of Datasets
(Details in Appendix: A.5, Table: 8). Summarization: We use the popular CNN-DailyMail (CNN/DM) (Hermann et al., 2015) dataset which is a large collection (over 300k) of text summarization pairs, created from CNN and Daily Mail news articles. Each datapoint in this dataset comprises of an article (the body of the news article with 683 words on average) and the corresponding highlights (article summary as written by the article author). While the training set contains more than 287k samples, we have randomly chosen 4k samples for training both routers and LoRA. During training in our framework, the number of trainable parameters is small in both phases - therefore a small subset of data points is sufficient for training.
Inference is performed on the standard test set with 11,490 samples.
Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.0) (Rajpurkar et al., 2016), a widely-used benchmark for machine Question Answering. The dataset consists of over 100k question-answer pairs posed by crowd-workers on a set of over 500 Wikipedia articles. Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage) and the corresponding answer (a text span from the corresponding reading passage). Similarly to the CNN/DM dataset, 4k samples are chosen as random to train both routers and LoRA. The training and validation split contains 87,599 and 10,570 samples respectively. Evaluation is performed on the validation set Schuster et al. (2022) as the test set labels are not publicly released.
@@ A.3 Evaluation Metrics
  • BLEU Score: BLEU (Bilingual Evaluation Understudy) scores are used to measure the quality of translations. BLEU compares n-grams of the candidate translation to n-grams of the reference translation, providing a score between 0 and 1, with higher scores indicating better translations. In this evaluation, NLTK BLEU is employed, focusing on BLEU-1 and BLEU-2 scores.
@@ A.4 Training and Inference Setup
  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively. Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization. The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps. The learning rate is set between $1e^{-4}$ and $3e^{-4}$ a cosine scheduler is used to adjust the learning rate. Gradients are accumulated after 5 steps and the regularization coefficient $\lambda$ is fixed at 0.01. For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor (lora_alpha) of 32. We set the non-skipping penalization regularizer hyper-parameter $\beta=\alpha/3$ after tuning to maintain skipping level - where $\alpha$ is the non-skipping penalization hyperparameter used in the first phase and set according to the target speedup ratio. For translation, the maximum sequence length is set to 128 for router training and 256 for LoRA training. Similarly, for summarization, the maximum sequence length is set to 500 and 700 respectively. For Question Answering, this length is set to 512. Prompts for the different tasks regarding training/inference are shown in Appendix A.6.
  • Inference settings: For all the tasks, we set the temperature to 0.8 and enable top-k sampling over 10 tokens. The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD. Caching is turned on during inference.
A.5 Training and testing split
WMT Summarization RC
Eng-to-German Eng-to-Chinese CNN/DM SQuAD
Train 3505 8983 3400 3400
Validation 876 998 600 600
Test 2038 2038 11490 10570
Table 8: Train-Validation-Test split for WMT, CNN/DM and SQuAD datasets
@@ A.5 Training and testing split
A.6 Prompt Details
@@
  • For the Question Answering task (used in SQuAD dataset), the following prompt structure is utilized:
@@
A.7 Layer-wise Skipping Statistics
Table 15 present the fraction of sequences that skip a particular block during the task for the LLaMA-3-8B model. If the corresponding cell in a row shows a value of 80.00, it implies that 80% of the sequences skip this block. It is important to note that the decision regarding which block to skip varies across different datasets and tasks. Additionally, partial skipping in some blocks, with varying percentages, suggests that while some sequences consider the layer important, others do not and therefore skip it during the decoding phase.
Figures 3, 4, 5, and 6 depict the distribution of skipped blocks when the model is configured to skip approximately 15% of the layers. These plots demonstrate the variation in block skipping based on the task and dataset, highlighting the differences in block importance. The bar plots further show that the determination of important layers depends not only on the dataset but also on the model architecture. For instance, Figure 3 illustrates that layers 7 and 9 are entirely skipped in LLaMA-3-8b, while they are rarely skipped in LLaMA-3.2-3b.
A.8 Detailed Result Table
Tables 9 and 11 present the detailed results for the LLaMA-3.2-3B model, while Tables 10 and 12 summarize the performance of the LLaMA-3-8B model for an additional skipping percentage on Machine Translation Task. The results are reported using BLEU (BLEU-1, BLEU-2) and COMET metrics, highlighting performance across different skipping percentages. Similarly, Table 13 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L. Lastly, Table 14 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72
Base 31.19 13.67 81.66
15 Skip Decode Router + LoRA 23.24 11.07 44.58
Router 9.93 3.89 32.74
Random Skip Router + LoRA 18.99 4.78 47.26
Router 12.11 2.77 36.30
Unified Skip Router + LoRA 21.65 5.93 44.72
Router 16.41 3.99 39.81
FiRST (Ours) Router + LoRA 24.68 7.13 60.29
Router 16.47 4.11 43.04
25 Skip Decode Router + LoRA 16.85 7.62 32.33
Router 9.14 3.51 27.64
Random Skip Router + LoRA 8.95 1.07 30.97
Router 5.32 0.69 27.22
Unified Skip Router + LoRA 15.14 2.65 39.81
Router 10.21 1.62 30.86
FiRST (Ours) Router + LoRA 18.89 4.06 45.38
Router 9.51 1.44 29.08
35 Skip Decode Router + LoRA 1.62 0.36 23.30
Router 1.71 0.27 19.34
Random Skip Router + LoRA 4.72 0.15 25.08
Router 1.96 0.06 21.18
Unified Skip Router + LoRA 1.18 0.05 19.65
Router 0.90 0.03 20.36
FiRST (Ours) Router + LoRA 9.47 1.29 27.45
Router 5.74 0.56 25.03

### v3

@@ 5.1 Baselines for comparison
  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method relies on using a heuristic-based strategy for retaining layers at fixed intervals. We replicate the algorithm in (Liu et al., 2024) and compare performance both with and without LoRA fine-tuning across various skipping percentages.
@@ 5.2 Experimental Results on Different Tasks
WMT: For LLaMA-3-8B, at $15\%$ skipping, FiRST achieves a latency improvement of upto 10% on TPOT (see Tables 1 and 4).In most cases, it significantly outperforms other layer skipping strategies (Skip Decode, Random and Unified Skipping) and in other cases, it is comparable in quality. When compared to the gold output (Base + LoRA), FiRST generally retains a high percentage of the quality, for instance, often achieving $\geq 80\%$ of the COMET score for LLaMA-3-8B, and $\geq 70\%$ in BLEU scores. For $25\%$ skipping, FiRST achieves significant improvement in quality over other strategies, in almost all metrics, while achieving $\sim 18\%$ reduction in TPOT.
For LLaMA-3.2-3B, for similar latency improvement ($\sim 10\%$), the COMET scores are significantly higher that other baselines while the BLEU scores are comparable. (Table 4). It remains within $65-85\%$ of BLEU and COMET scores (Table 1) achieved by the gold standard.
@@
CNN/DailyMail Dataset: For LLaMA-3-8B, at $\sim 15\%$ skipping, our method outperforms the Base + LoRA setting(Table: 3) while obtaining a 12% improvement in TPOT (Table 4). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
Model size Model Type WMT CNN/DM SQuAD Natural Questions
Skip (%) Eng→De Eng→Zh Skip (%) TPOT Skip (%) TPOT Skip (%) TPOT
LLaMA-3–8.B Base + LoRA 0 0 0 0
R + LoRA 15 0.90× 0.88× 15 0.88× 10 0.95× 10 0.96×
R + LoRA 25 0.82× 0.83× 20 0.81× 20 0.78× 20 0.80×
LLaMA-3.2–3B Base + LoRA 0 0 0 0
R + LoRA 15 0.90× 0.91× 15 0.88× 15 0.94× 15 0.94×
R + LoRA 25 0.78× 0.75× 24 0.79× 24 0.83× 24 0.84×
@@
Table 4: TPOT variation across all datasets for FiRST. The reported values are relative to the LoRA fine-tuned base model. R + LoRA indicates Router Augmentation followed by LoRA fine-tuning. Fine-tuning improves TPOT and quality significantly.
@@
SQuAD Dataset: For the LLaMA-3-8B model, FiRST(at 10% skip level) maintains over 95% of the performance of the gold standard (Base + LoRA without skipping)(Table 2), with overall latency gains upto 16% (Table 4). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics.
Natural Questions Dataset: For the LLaMA-3-8B model, FiRST retains over $92\%$ of the EM achieved by the non-skipping gold standard (Table 2), with overall latency gains of 5-12% (Table 4). It is significantly better in quality than all other baselines for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies. While more sophisticated retrieval strategies could further enhance performance, our goal was to demonstrate that FiRST maintains quality performance and latency gains even in a retrieval-based QA setting.
Detailed results for an additional skipping percentage are provided in Appendix A.7.
@@
Layer-wise Skipping Patterns: Layer-wise skipping varies significantly across tasks, reflecting the task-specific importance of each layer. For LLaMA-3-8B at a 15% skipping rate, layers 7–9 and 21 are fully skipped in English-to-German, with partial skipping in layer 18. Figure 3 for summarization shows that only some layers in the middle of the network, specifically layers 19 to 27, are skipped. The early layers and the last few layers (28–32) are never skipped. Layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19 and 21. Some layers are skipped less than 10% of times, indicating their necessity for specific sequences. This shows that it’s important to learn which layers to skip based on each input, as the skipping pattern is not the same for every sample. The task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped on SQuAD while layers 12, 23 are fully skipped in case of NQ dataset, and skipping patterns depend on the input. Detailed statistics are in Appendix A.6. Furthermore, the router is trained in a task-aware manner, ensuring that skipping decisions align with task complexity. We also conduct experiments to validate the reusability and generalizability of routers (see Appendix: A.8) across different datasets for the same underlying task.
Figure 3: LLaMA-3-8B skipping statistics at 15% skipping rate on summarization task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Computational overhead of Routers: We aim to improve the efficiency of pre-trained models by introducing routers, lightweight linear classifiers that help decide whether to skip certain layers during inference. A common concern with adding such components is the potential increase in computational cost. However, our analysis shows that the additional parameters and operations introduced by the routers account for a very small fraction of the total model computation, only 0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B in terms of parameter count, and around 0.3% of the total forward pass time. This fixed cost remains negligible compared to the overall model execution time. As a result, the overall latency gain primarily depends on how much computation is skipped, rather than the routing mechanism itself. Given similar levels of skipping, different methods tend to show comparable efficiency gains, so the key focus becomes minimizing the performance drop relative to the original model. Our results show that the proposed routing mechanism achieves this balance effectively, preserving latency improvements while maintaining model quality.
@@ 6 Conclusion
We propose a new framework, FiRST, for layer selection that adapts to the input sequence and task, aiming to reduce latency in a quality-aware manner. This approach is sequence-dependent and operates compatibly with KV caching. With an optimal layer skipping rate of around 15%, FiRST achieves a 10-20% reduction in latency. This speedup is achieved while remaining quality-neutral, maintaining approximately 80% or more of the base model’s performance on quality metrics across multiple tasks (Machine Translation, Summarization, Question Answering) and model architectures on well-known open-source datasets. Furthermore, our method significantly outperforms other layer selection strategies on most quality metrics.
@@ A.1 Related Work
Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022a; Wu et al., 2020) where adaptive compute is used for different parts of the token sequence. While these approaches have been popular for encoder-only models which processes the entire sequence as a whole, they have faced challenges in generation tasks. The main limitation of these set of techniques are their inability to handle KV caching appropriately which is crucial for multi-fold speed-ups in current LLM architectures. We emphasize that in our work, we assign varying compute to sequences in different batches but within the same sequence, we assign the same compute to every token.
@@
Mixture of Experts
Mixture of Experts (MoE) is a well-established technique for improving the efficiency and capacity of deep learning models by conditionally activating subsets of parameters for different inputs. MoE-based transformer models, such as Switch Transformers (Fedus et al., 2021) and GShard (Lepikhin et al., 2020), employ a gating mechanism to route tokens to a subset of expert layers, thereby significantly reducing computational costs while maintaining expressivity. These architectures are designed to scale up model capacity without a proportional increase in inference cost. The main goal is to scale up the parameters while maintaining the cost of pre-training and inference.
Although MoE and input-adaptive layer skipping share the goal of selective computation, they differ in fundamental ways. MoE dynamically routes tokens to different experts at the layer level, whereas layer skipping focuses on bypassing entire layers in the transformer stack based on the input query. MoE models typically maintain a full-depth model structure, leveraging sparse activation to reduce computational overhead, whereas layer-skipping methods explicitly modify the depth of computation for different inputs. Simply put, the techniques are orthogonal - layer skipping can be used in tandem with Moe to further reduce the depth of computation wherever possible at an expert level
@@ A.2 Details of Datasets
WMT Summarization Question Answering
[…]
Train 3,505 8,983 3,400 3,400 3,400
Validation 876 998 600 600 600
Test 2,038 2,038 11,490 10,570 7,830
Table 5: Train, validation, and test splits for machine translation (WMT), summarization (CNN/DM), and question answering (SQuAD, NQ) tasks.
@@
Summarization: We use the popular CNN-DailyMail (CNN/DM) (Hermann et al., 2015) dataset which is a large collection (over 300k) of text summarization pairs, created from CNN and Daily Mail news articles. Each datapoint in this dataset comprises of an article (the body of the news article with 683 words on average) and the corresponding highlights (article summary as written by the article author). While the training set contains more than 287k samples, we have randomly chosen 4k samples for training both routers and LoRA. During training in our framework, the number of trainable parameters is small in both phases - therefore a small subset of data points is sufficient for training.
Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.1) (Rajpurkar et al., 2016), a widely-used benchmark for Machine Question Answering. The dataset consists of over 100k question-answer pairs posed by crowd-workers on a set of over 500 Wikipedia articles. Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage), and the corresponding answer(a text span from the corresponding reading passage). Similarly to the CNN/DM dataset, 4k samples are chosen at random to train both routers and LoRA. The training and validation splits contain 87,599 and 10,570 samples, respectively. Evaluation is performed on the validation set Schuster et al. (2022a) as the test set labels are not publicly released. Natural Questions (NQ): We use the Natural Questions dataset (Kwiatkowski et al., 2019), which consists of real queries issued to Google Search paired with relevant Wikipedia articles. Each example in NQ contains a query (an actual user question), a context(Wikipedia article), and two types of answers: a long answer (typically a paragraph) and a short answer (a specific text span, when available). We implement a naïve RAG solution using a simple S-BERT model to generate embeddings for the query and passages and retrieve the top 5 most relevant passages based on similarity. Once retrieved, these passages are used as context to answer the query and then compared against gold answers. Similar to SQuAD, 4k samples are randomly chosen to train both routers and LoRA. The training set contains over 300k examples, with evaluation performed on 7.83k validation samples containing short answers. For further details on training-testing split, refer to table: 5
@@ A.3 Evaluation Metrics
  • BLEU Score: BLEU (Bilingual Evaluation Understudy) scores are used to measure the quality of translations. BLEU compares n-grams of the candidate translation to n-grams of the reference translation, providing a score between 0 and 1, with higher scores indicating better translations. In this evaluation, NLTK BLEU is employed and We report BLEU-1, BLEU-2, and the cumulative BLEU score, which is computed as the geometric mean of individual n-gram precision scores from unigram to 4-gram. To mitigate the issue of zero counts for higher-order n-grams, we apply the smoothing strategy utilized in Post (2018).
@@ A.4 Training and Inference Setup
[…]
  • Joint training considerations: The decision to split training into two phases—first training the router and then the LoRA modules—was made to ensure better stability during optimization. In our experiments, this approach allowed the router to establish a robust skipping strategy without interference from the fine-tuning of LoRA modules, which otherwise introduces instability. While training both the router and LoRA modules together is theoretically possible and could eliminate the need for reapplying the non-skip penalty in the second phase, initial experiments revealed that it led to suboptimal convergence and conflicting gradients. This was particularly evident when the router’s decisions were not yet stable, causing inconsistencies in the joint optimization process.
  • Inference settings: For all the tasks, we set the temperature to 0.8 and enable top-k sampling over 10 tokens. The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD and NQ. Caching is turned on during inference.
@@
Figure 8: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 10% skipping rate on Natural Questions Answering Task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
A.5 Prompt Details
@@
  • For the Question Answering task (used in SQuAD/NQ dataset), the following prompt structure is utilized:
@@
A.6 Layer-wise Skipping Statistics
To illustrate broader layer-wise behaviors, Figures 4 through 8 show block-wise skip rates observed when the models were configured for an average skip rate of approximately 15%. These figures highlight how individual layers exhibit distinct skipping behaviors across various tasks, including translation (English-German, English-Chinese), summarization (CNN/DM), and question answering (SQuAD, Natural Questions). It also reveals that tasks with similar processing requirements (e.g., different translation language pairs) exhibit comparable skipping patterns. Providing a more granular view for specific models, Table 8 reports the fraction of sequences that skip each block under LoRA adaptation for the LLaMA-3-8B and LLaMA-3.2-3B models. These fractions are presented as decimals, so a value of 0.10 signifies that 10% of sequences skip the corresponding block.
Skip (%) Model Type Config BLEU-1 BLEU-2 BLEU COMET
LLaMA-3.2-3B
0 Original Model Base + LoRA 0.376 0.174 0.160 89.72
Base 0.312 0.137 0.125 81.66
15 Skip Decode Router + LoRA 0.207 0.069 0.105 44.58
Router 0.099 0.039 0.055 32.74
Random Skip Router + LoRA 0.190 0.048 0.079 47.26
Router 0.121 0.028 0.048 36.30
Unified Skip Router + LoRA 0.216 0.059 0.095 44.72
Router 0.164 0.040 0.067 39.81
FiRST (Ours) Router + LoRA 0.247 0.071 0.113 60.29
Router 0.165 0.041 0.069 43.04
25 Skip Decode Router + LoRA 0.168 0.047 0.077 32.33
Router 0.091 0.035 0.050 27.64
Random Skip Router + LoRA 0.090 0.011 0.023 30.97
Router 0.053 0.007 0.015 27.22
Unified Skip Router + LoRA 0.151 0.027 0.050 39.81
Router 0.102 0.016 0.032 30.86
FiRST (Ours) Router + LoRA 0.189 0.041 0.072 45.38
Router 0.095 0.014 0.029 29.08
35 Skip Decode Router + LoRA 0.058 0.011 0.018 23.30
Router 0.017 0.003 0.005 19.34
Random Skip Router + LoRA 0.047 0.001 0.007 25.08
Router 0.020 0.001 0.003 21.18
Unified Skip Router + LoRA 0.012 0.000 0.002 19.65
Router 0.009 0.000 0.001 20.36
FiRST (Ours) Router + LoRA 0.095 0.013 0.027 27.45
Router 0.057 0.006 0.014 25.03
LLaMA-3-8B
0 Original Model Base + LoRA 0.418 0.217 0.199 93.00
Base 0.372 0.186 0.169 87.13
15 Skip Decode Router + LoRA 0.230 0.105 0.094 55.62
Router 0.040 0.012 0.019 23.33
Random Skip Router + LoRA 0.304 0.110 0.097 66.25
Router 0.265 0.088 0.132 60.27
Unified Skip Router + LoRA 0.289 0.106 0.153 59.34
Router 0.232 0.079 0.117 59.26
FiRST (Ours) Router + LoRA 0.380 0.179 0.161 82.14
Router 0.288 0.118 0.108 67.74
25 Skip Decode Router + LoRA 0.118 0.052 0.070 31.47
Router 0.032 0.009 0.015 21.55
Random Skip Router + LoRA 0.060 0.009 0.018 29.71
Router 0.037 0.005 0.011 29.95
Unified Skip Router + LoRA 0.157 0.034 0.060 31.69
Router 0.126 0.027 0.048 32.15
FiRST (Ours) Router + LoRA 0.178 0.041 0.071 34.95
Router 0.097 0.014 0.029 26.01
35 Skip Decode Router + LoRA 0.049 0.020 0.028 23.85
Router 0.030 0.008 0.013 20.03
Random Skip Router + LoRA 0.018 0.001 0.004 25.56
Router 0.014 0.001 0.002 25.34
Unified Skip Router + LoRA 0.064 0.008 0.017 22.05
Router 0.039 0.005 0.011 22.88
FiRST (Ours) Router + LoRA 0.064 0.004 0.012 19.96
Router 0.037 0.001 0.005 21.41